### v1

# Directed and Undirected Vertex Connectivity Problems are Equivalent for Dense Graphs

Abstract
Vertex connectivity and its variants are among the most fundamental problems in graph theory, with decades of extensive study and numerous algorithmic advances. The directed variants of vertex connectivity are usually solved by manually extending fast algorithms for undirected graphs, which has required considerable effort.
In this paper, we present a simple, black-box randomized reduction from directed to undirected vertex connectivity for dense graphs. As immediate corollaries, we largely simplify the proof for directed vertex connectivity in $n^{2+o(1)}$ time [LNP+25], and obtain a parallel vertex connectivity algorithm for directed graphs with $n^{\omega+o(1)}$ work and $n^{o(1)}$ depth, via the undirected vertex connectivity algorithm of [BJMY25]. Our reduction further extends to the weighted version of the problem. By combining our reduction with the recent subcubic-time algorithm for undirected weighted vertex cuts [CT25], we obtain the first subcubic-time algorithm for weighted directed vertex connectivity, improving upon a three-decade-old bound [HRG00] for dense graphs.
1 Introduction
Vertex connectivity and its variances are among the most fundamental graph problems and have received extensive studies over the last several decades [Kle69, Pod73, ET75, Eve75, Gal80, EH84, Mat87, BDD+82, LLW88, CT91, NI92, CR94, Hen97, HRG00, Gab06, CGK14, NSY19, FNY+20, SY22, NSY23, BJMY25, Kor25, CT25, JNSY25]. For an undirected graph, a vertex cut is a vertex set whose removal disconnects the graph. The problem of vertex connectivity asks to find a minimum vertex cut of a non-complete undirected graph. For directed graphs, a vertex cut is a vertex set whose removal renders the graph not strongly connected. The problem of directed vertex connectivity is to find a minimum vertex cut on a non-complete directed graph.
For a long time, progress in solving directed vertex connectivity often involved a manual, ad-hoc process of “mimicking” fast algorithms from the undirected graphs. For example, consider the fastest known algorithm [^1] Unless stated otherwise, all algorithms in this paper are randomized Monte Carlo algorithms. A gap remains between randomized and deterministic algorithms for vertex connectivity. We refer to [JNSY25] for more discussion. for vertex connectivity. A long line of research [NSY19, FNY+20, LNP+25] culminates in an almost linear-time randomized algorithm for undirected vertex connectivity. For directed graphs, [LNP+25] extends their techniques and spends 9 more pages to obtain an $n^{2+o(1)}$-time randomized algorithm for directed vertex connectivity, where $n$ denotes the number of vertices.
A similar phenomenon appeared in the algebraic approach, which was first proposed for undirected graphs [LLW88]; later, the same approach was manually adapted to work for directed graphs [CT91]. Very recently, this approach has been made faster on undirected graphs using common-neighborhood clustering techniques to obtain an $n^{\omega+o(1)}$-time algorithm [BJMY25] where $\omega\leq 2.371339$ is the matrix multiplication exponent [ADV+25]. It is conceivable that their techniques can also be extended to directed graphs, but this has not been done yet.
This suggested a deep connection between undirected and directed problems. In this paper, we resolve the missing link by introducing a simple black-box randomized reduction from directed to undirected vertex connectivity for dense graphs. This formalizes the long-observed phenomenon that the problems were related and explains why mimicking was so often successful.
Theorem 1.1 (Informal version of Theorem 3.1).
The directed vertex connectivity problem on an $n$-vertex directed graph reduces to undirected vertex connectivity problems with $O(n^{2}\log n)$ edges.
The reduction is short and elementary, which is surprising given that vertex connectivity has received considerable attention over the past five decades.
As a corollary of Theorem 1.1, we obtain the following results immediately. Combining with $m^{1+o(1)}$ undirected vertex connectivity algorithm by [LNP+25], we obtain:
Corollary 1.2.
There is an $n^{2+o(1)}$-time directed vertex connectivity algorithm.
Thus, it is no longer necessary to open the box and mimic the proof of undirected vertex connectivity as done in [LNP+25]. Furthermore, since the reduction is fully parallelizable, we also obtain parallel algorithms in this setting. Combining with $n^{\omega+o(1)}$-time undirected vertex connectivity algorithm by [BJMY25], we obtain:
Corollary 1.3.
There is a directed vertex connectivity algorithm with $n^{\omega+o(1)}$-work and $n^{o(1)}$-parallel depth.
Work refers to the total amount of computation across all processors. Parallel depth refers to the length of the longest chain of dependent operations in an algorithm.
(Weighted) Vertex Mincut.
[…]
Interestingly, our reduction Theorem 1.1 also holds on weighted graphs. Strictly speaking, we have the following result.
Theorem 1.4 (Informal version of Theorem 3.1).
The directed minimum vertex cut problem on an $n$-vertex directed graph reduces to the undirected minimum vertex cut problem with $O(n^{2}\log n)$ edges.
By plugging in Theorem 1.4 with the $\min(mn^{0.99+o(1)},m^{1.5+o(1)})$ undirected minimum vertex cut algorithm [CT25], we immediately get the following corollary that improves the $\widetilde{O}\left(mn\right)$ algorithm [HRG00] from 30 years ago on dense graphs.
Corollary 1.5.
There is an $n^{2.99+o(1)}$ time randomized algorithm for directed minimum vertex cut.
Beyond the sequential model.
We believe the simplicity of our reduction will make it easy to extend to other computation models, including communication, distributed, or streaming models [BJMY25]. These models are especially suitable for our reduction theorem because many of them use $n$ as the parameter. However, individual implementation involves delicate technical details of the reduction that are beyond the scope of this paper.
Comparison to Edge Connectivity.
Edge connectivity asks for the minimum number of edges whose removal disconnects an undirected graph or breaks the strong connectivity of a directed graph. Although this problem appears very similar to vertex connectivity, there is no known reduction between undirected and directed edge connectivity, making our result even more surprising. In fact, directed edge connectivity is generally considered harder to handle than its undirected counterpart. It has long been known that undirected edge connectivity can be solved in nearly linear time using (relatively) simple combinatorial techniques [Kar00]. In contrast, directed edge connectivity implies solving unit-capacity maximum flow, which was only recently shown to be solvable in almost linear time even on dense graphs [vdBLL+21, CKL+25], using sophisticated continuous optimization methods combined with complex data structures.
Concluding Remarks.
One can notice that our reduction from directed to undirected graph has a loss in boosting the number of edges to $n^{2}$, which is not a good reduction when the graph is moderately sparse. This also collides with the fact that directed vertex connectivity has remained at $n^{2+o(1)}$ running time for some time now. It is a great open problem to prove that directed vertex connectivity can be solved in almost linear time for any sparsity, not necessarily done by a black-box reduction. Another open problem is whether the reduction can be made deterministic. This seems plausible since vertex connectivity can be solved deterministically in $m^{1+o(1)}k$ time on unweighted undirected graphs where $k$ is the vertex connectivity, and $\tilde{O}(m^{1+o(1)}n)$ time on weighted directed graphs [JNSY25].
2 Preliminaries
Graphs.
In this paper, we consider both directed and undirected graphs $G=(V,E)$. $E\subseteq V\times V$ is the edge set, and when $G$ is an undirected graph, we consider $(u,v)$ and $(v,u)$ as the same edge. When we say $G$ is vertex-weighted, each vertex is assigned a weight $w_{G}(v)\in{\mathbb{Z}}^{+}$ (when $G$ is clear from the context, we omit $G$ and write $w(v)$). We always assume $w(v)$ is a polynomially bounded integer for simplicity. We write $G^{T}$ as the graph obtained by reversing the arc of $G$.
We use $N_{G}(u)\subseteq V$ to denote the neighborhood set of $u$ on $G$ when $G$ is an undirected graph. When $G$ is directed, we use $N^{out}_{G},N^{in}_{G}$ to denote the out and in neighbors. When we use $N^{out}_{G},N^{in}_{G}$ on an undirected graph, both of them refer to $N_{G}$. We write $N_{G}[u]=N_{G}(u)\cup\{u\}$, similar for other notations. When $G$ is clear from the context, we omit $G$ from the subscript. The minimum degree of a graph $G$ is defined as $\min_{v\in V}|N(v)|$, and the minimum weighted-degree is defined as $\min_{v\in V}(\sum_{u\in N(v)}w(u))$.
Vertex cuts.
A vertex cut $(L,S,R)$ is a tri-partition of the vertex set $V$ such that there are no edges from $L$ to $R$ where both $L$ and $R$ are non-empty sets. The vertex set $S$ is called a separator (following the literature, we also refer to it as a vertex cut). The size of the vertex cut is $|S|$ and the weight of the vertex cut is $w(S):=\sum_{v\in S}w(v)$. If $s\in L,t\in R$, then we call $(L,S,R)$ (or $S$) a $(s,t)$-vertex cut. It is an $s$-vertex cut if $s\in L$. Throughout this paper, we assume that every input graph is non-complete and thus a vertex cut always exists.
Minimum vertex cut.
The problem of minimum vertex cut asks to find a vertex cut $S$ that minimizes the weight $w(S)$ (in the case of a weighted graph) or the size $|S|$ (in the case of an unweighted graph). The latter case is also referred to as vertex connectivity.
Single-source minimum vertex cut.
The problem of single source minimum vertex cut is given a graph and a source $s\in V$, asks to find an $s$-vertex cut $(L,S,R)$ that minimizes $w(S)$ (in the case of a weighted graph) or the size $|S|$ (in the case of an unweighted graph). We denote $\kappa_{G}(s)=w(S)$ or $|S|$ in the case of an unweighted graph.
3 A Reduction from Directed to Undirected Vertex Connectivity
We state our main result.
Theorem 3.1.
Suppose there is an algorithm ${\mathcal{A}}_{0}$ solving the minimum vertex cut on undirected graphs. In that case, there is a randomized algorithm solving minimum vertex cut on directed graphs w.h.p., making calls to ${\mathcal{A}}_{0}$ with cumulative $O(n^{2}\log n)$ many edges and additional $O(n^{2}\log n)$ time. Moreover, the reduction holds even when we restrict both inputs to unweighted graphs.
To prove Theorem 3.1, the reduction can be done in three steps as follows.
  1. 1.
    We reduce directed vertex connectivity (Directed VC) to directed single-source vertex connectivity (Directed SSVC) in Section 3.1.
  2. 2.
    Then, we reduce directed single-source vertex connectivity (Directed SSVC) to undirected single-source vertex connectivity (Undirected SSVC) in Section 3.2.
  3. 3.
    Finally, we reduce undirected single-source vertex connectivity (Undirected SSVC) to undirected vertex connectivity (Undirected VC) in Section 3.3.
The constructions in Sections 3.1 and 3.3 are folklore. The key novelty is the reduction in Section 3.2.
3.1 A Reduction from Directed VC to Directed SSVC
Lemma 3.2.
If there is an algorithm ${\mathcal{A}}_{0}$ solving single-source minimum vertex cut on directed graphs, then there is a randomized algorithm solving minimum vertex cut on directed graphs w.h.p., making calls to ${\mathcal{A}}_{0}$ with cumulative inputs $O(m\log n)$ edges and additional $O(n^{2})$ time. Moreover, the reduction holds even when we restrict both input graphs to be unweighted or undirected.
Proof.
We start by describing the algorithm. We assume the graph is not a complete graph as otherwise there is no vertex cut. Let the minimum weighted-degree of the input graph be $\delta$, and let $W=\sum_{v\in V}w(v)$. The algorithm samples $O(\log n\cdot\frac{W}{W-\delta})$ many nodes independently, where each node $u$ get sampled with probability $w(u)/W$. For each sampled node $u$, define $G_{u}$ as the graph obtained from deleting every edge $(v_{1},v_{2})$ with $v_{1},v_{2}\in N_{G}(u)$ from $G$. The algorithm calls ${\mathcal{A}}_{0}$ on $G_{u}$ with source $u$ for every sampled node $u$, and returns the smallest vertex cut among all returns. We repeat the same algorithm for $G^{T}$ and take the smaller vertex cut returned.
Correctness.
Let $(L,S,R)$ be one of the smallest vertex cut. Suppose without loss of generality that $w(L)\geq w(R)$ (one of $G$ or $G^{T}$ must satisfy this condition). Notice that $\delta\geq w(S)$ because deleting vertices of weight $\delta$ disconnects the graph, and $S$ should be the smallest weight vertex set whose deletion disconnect the graph. Thus, we get $w(L)\geq(W-\delta)/2$. According to Chernoff’s bound, w.h.p. there is a sampled node $u\in L$, whose single-source minimum vertex cut has weight at most $w(S)$ since $S$ is clearly a single-source vertex cut in $G_{u}$ (since we only delete edges from $G$ to get $G_{u}$). To see that the returned cut in $G_{u}$ must be a cut in $G$, let $(L^{\prime},S^{\prime},R^{\prime})$ be a vertex cut in $G_{u}$ with $u\in L^{\prime}$. Then we have $N_{G}(u)\subseteq L^{\prime}\cup S^{\prime}$ since all edges adjacent to $u$ is present in $G_{u}$. Thus, we only delete edges between $L^{\prime}\cup S^{\prime}$ from $G$ to get $G_{u}$, which means $(L^{\prime},S^{\prime},R^{\prime})$ is still a cut in $G$.
Complexity.
We will analyze the expected size of the summation of all the graphs input to ${\mathcal{A}}_{0}$. Let $u$ be a sampled node. Fix an edge $e\in E$, let $I_{e}$ be the event that $e\in G_{u}$. Notice that if $I_{e}$ happens, then one of $e$’s end point must not be a neighbor of $u$. In other words, let $e=(v_{1},v_{2})$, then $u$ must be in either $V-N^{in}_{G}(v_{1})$ or $V-N^{in}_{G}(v_{2})$. Moreover, we have $w(V-N^{in}_{G}(v_{1})),w(V-N^{in}_{G}(v_{2}))\leq W-\delta$. Thus, the probability that $I_{u}$ happens is at most $2(W-\delta)/W$. Notice that the size of $G_{u}$ is at most $\sum_{e\in E}I_{e}$, which means the expectation of the size of $G_{u}$ can be upper bounded by $m\cdot 2(W-\delta)/W$. Recall that we sampled $O(\log n\cdot\frac{W}{W-\delta})$ many nodes, so the total number of edges among all $G_{u}$ has expectation $O(m\log n)$. Notice that building all the graphs $G_{u}$ can be done by brute force in $O(n^{2})$ time.
Notice that when $G$ is unweighted or undirected, the graph $G_{u}$ becomes unweighted or undirected as well. ∎
3.2 A Reduction from Directed SSVC to Undirected SSVC
Lemma 3.3.
Given an algorithm ${\mathcal{A}}_{0}$ solving single-source minimum vertex cut on undirected graphs, there is an algorithm that solves the single-source minimum vertex cut problem on directed graphs, making one call to ${\mathcal{A}}_{0}$ with input of $O(n^{2})$ edges and requiring additional $O(n^{2})$ time. Moreover, the reduction holds even when we restrict both inputs to unweighted graphs.
We first describe the construction and its property.
Construction.
We are given a weighted directed graph $G=(V,E,w)$ and a source $s$. Define $A=V-N^{out}_{G}[s]$ and $\forall x\in A,w^{\prime}(x)=w(x)$, and define $A^{\prime}=\{u^{\prime}\mid u\in A\}$ as a copy of $A$ with the same node weights, i.e., $\forall x^{\prime}\in A^{\prime},w^{\prime}(x^{\prime})=w(x)$.
We define the skeleton $H=(V^{\prime},E_{H},w^{\prime})$ where
$V^{\prime}$ $=V\cup A^{\prime}$
$E_{H}$ $=E_{1}\cup E_{2}\cup E_{3}\mbox{ where }$
$E_{1}$ $=\{(s,v)\mid v\in A^{\prime}\cup N^{out}_{G}(s)\},\mbox{ connecting }s\mbox{ to all vertices but }A,$
$E_{2}$ $=\{(u,u^{\prime})\mid u\in A\},\mbox{a perfect matching between }A\mbox{ and }A^{\prime}$
$E_{3}$ $=\{(u,v)\mid u,v\in A\},\mbox{a clique inside }A.$
Note that the skeleton depends on the fixed set $N^{out}_{G}(s)$. To complete the construction, we add the remaining edges from $G$ to $H$.
We define a weighted undirected graph $G^{\prime}=(V^{\prime},E^{\prime}):=H\cup E_{4}\cup E_{5}$ where
$E_{4}=$ $\{(u,v^{\prime})\mid u,v\in A,(v,u)\in E\},$
$E_{5}=$ $\{(u,v)\mid u\in A,v\in N^{out}_{G}(s),(v,u)\in E\}.$
In $E_{4}$, for every edge $(v,u)\in E$ where $u,v\in A$, we lift its tail to $A^{\prime}$, i.e., we add $(u,v^{\prime})$ to $E_{4}$. In $E_{5}$, for every edge $(v,u)\in E$ going out of $N_{G}^{out}(s)$, we add $(u,v)$ to $E_{5}$. We define $F_{4}$ and $F_{5}$ as the set of edges in $G$ that are mapped to $E_{4}$ and $E_{5}$, respectively. This completes the construction, see Figure 1 for illustration.
Figure 1: (Left) The skeleton graph $H$. (Right) The complete construction. The set $R\subseteq A$ is denoted as a red box.
Before we describe the reduction, let us observe useful properties of $G^{\prime}$. The role of $E_{1}$ is to ensure the following proposition.
Proposition 3.4.
For every vertex cut $(L^{\prime},S^{\prime},R^{\prime})$ in $G^{\prime}$ if $s\in L^{\prime}$ then $R^{\prime}\subseteq A$
Proof.
Suppose $R^{\prime}$ contains some node $x\not\in A$. Thus, $s$ does not have an edge to $x$ by the definition of a vertex cut. However, by $E_{1}$, $s$ has an edge to all but $A$. In particular, $s$ has an edge to $x$, a contradiction.∎
The crux of the construction is to ensure the following property.
Lemma 3.5.
For all $\emptyset\neq T\subseteq A,$ $w(N_{G^{\prime}}(T))=w(N_{G}^{in}(T))+w(A).$
Proof.
The set $N_{G}^{in}(T)$ is equal to the following set. First, we start from $N_{G^{\prime}}(T)$ in $G^{\prime}$. Then, we drop $E_{2}\cup E_{3}$ from $G^{\prime}$ (removing all neighbors of $T$ from the skeleton $H$). The remaining neighbors of $T$ use $E_{4}\cup E_{5}$. In this case, we modify those edges in $E_{4}\cup E_{5}$ by undoing the lifting in $E_{4}$ and orienting the edges in $E_{5}$ appropriately, so that they correspond to the edges in $F_{4}\cup F_{5}$ in $G$. After this process, the total reduction in the weight of $N_{G^{\prime}}(T)$ is $w(N_{H}(T))=w(A)$. Hence, $w(N_{G}^{in}(T))=w(N_{G^{\prime}}(T))-w(N_{H}(T))$, which is equivalent to $w(N_{G^{\prime}}(T))=w(N_{G}^{in}(T))+w(N_{H}(T))=w(N_{G}^{in}(T))+w(A).$
This allows us to relate the $s$-vertex mincuts between $G$ and $G^{\prime}$.
Lemma 3.6.
$\kappa_{G^{\prime}}(s)=\kappa_{G}(s)+w(A)$.
Proof.
We break the proof into two directions.
  • $\kappa_{G^{\prime}}(s)\geq\kappa_{G}(s)+w(A)$. Let $(L^{\prime},S^{\prime},R^{\prime})$ be an $s$-vertex mincut in $G^{\prime}$. Since $s\in L^{\prime}$, Proposition 3.4 implies that $R^{\prime}\subseteq A$. Therefore,
    $\kappa_{G^{\prime}}(s)=w(N_{G^{\prime}}(R^{\prime}))=w(N_{G}^{in}(R^{\prime}))+w(A)\geq\kappa_{G}(s)+w(A)$
    The second equality follows from Lemma 3.5.
  • $\kappa_{G^{\prime}}(s)\leq\kappa_{G}(s)+w(A)$. Let $(L,S,R)$ be an $s$-vertex mincut in $G$. By design, $R\subseteq A$. Therefore,
    $\kappa_{G}(s)+w(A)=w(N_{G}^{in}(R))+w(A)=w(N_{G^{\prime}}(R))\geq\kappa_{G^{\prime}}(s)$
    The second equality follows from Lemma 3.5. ∎
We are ready to prove Lemma 3.3.
Proof of Lemma 3.3.
Given $G$ and $s$, we construct the aforementioned graph $G^{\prime}$. We call ${\mathcal{A}}_{0}$ on $G^{\prime}$ with source $s$. Suppose the output is $(L^{\prime},S^{\prime},R^{\prime})$ with $s\in L^{\prime}$. We return the vertex cut $N^{in}_{G}(R^{\prime})$. Note that $N^{in}_{G}(R^{\prime})$ is well-defined because $R^{\prime}\subseteq A$ by Proposition 3.4.
Correctness.
We prove that this is a vertex mincut in $G$ separating $s$ from $R$. Let $(L,S,R)$ be a vertex cut in $G$ where $R=R^{\prime},S=N^{in}_{G}(R^{\prime}),L=V(G)-(S\cup R)$. We first prove that $s\in L$. Indeed, by design, there is no edge from $s$ to $A$ in $G$. Since $R\subseteq A$, there is no edge from $s$ to $R$. It remains to prove $w(S)=\kappa_{G}(s).$ By Lemmas 3.5 and 3.6, $w(S)=w(S^{\prime})-w(A)=(\kappa_{G}(s)+w(A))-w(A)=\kappa_{G}(s).$
Time Complexity.
The graph $G^{\prime}$ can be constructed in $O(n^{2})$ time and the size of the graph is $O(n^{2})$. When $G$ is unweighted, $G^{\prime}$ is also unweighted. ∎
3.3 A Reduction from Undirected SSVC to Undirected VC
Lemma 3.7.
If there is an algorithm $A_{0}$ solving minimum vertex cut on undirected graphs, then there is an algorithm solving single-source minimum vertex cut on undirected graphs making one call to ${\mathcal{A}}_{0}$ with input $O(m)$ edges and additional $O(m)$ time. Moreover, the reduction holds even when we restrict both inputs to unweighted graphs.
Proof.
Given a single-source minimum vertex cut instance $G,s$, suppose $G$ is a weighted graph, we replace $s$ by a node with weight $W=\sum_{v\in V}w(v)$, then run ${\mathcal{A}}_{0}$ on $G$. The correctness and complexity are obvious, since $S$ cannot be in included in a minimum vertex cut.
If $G$ is an unweighted graph, we replace $s$ by a clique of size $\delta+1$ (the minimum degree of $G$) denoted by $C$, and connect all neighbors of $s$ to all vertices in the clique to get $G^{\prime}$. Let $(L,S,R)$ be a minimum vertex cut of $G^{\prime}$, firstly, $S$ cannot contain all vertices in $C$, as otherwise $S$ is larger than $\delta$. WLOG let $L$ be the vertex set that contain vertices in $C$. Then we have $C\cup N_{G}(s)\subseteq L\cup S$ according to the definition of $G^{\prime}$. Thus, we can remove all vertices in $C\cap S$ from $S$ to get $S^{\prime}$, and $S^{\prime}$ remains a vertex cut not containing $s$. In other words, if $S$ is minimum, then $S$ cannot contain any vertex in $C$. So calling ${\mathcal{A}}_{0}$ on $G^{\prime}$ suffices to find the single-source minimum vertex cut. The size of $G^{\prime}$ is $O(m)$ since there are at most $O(\delta\cdot n)$ newly added edges. ∎